OpenStudy (anonymous):

What is the probability of rolling a 6- sided die two time and getting a 3 and then a 2?

OpenStudy (perl):

P( 3 & 2) = P(3)* P(2) = 1/6 * 1/6

OpenStudy (mrnood):

the 2 events are independent therefore the prob of each is 1/6